OpenStudy (anonymous):

Identify the 16th term of a geometric sequence where a1 = 4 and a8 = -8,748. (2 points) Question 1 options: 1) -172,186,884 2) -57,395,628 3) 57,395,628 4) 172,186,884

OpenStudy (anonymous):

can you help me @shamim ?

OpenStudy (anonymous):

We can express a geometric sequence as : $$ \Large{ a_n = a_1\cdot r^{n-1}~,~~ n=1,2,3..} $$Now plug in the values given.

OpenStudy (anonymous):

plug in a1 = 4

OpenStudy (anonymous):

whats for r ?

OpenStudy (anonymous):

we can solve for r

OpenStudy (anonymous):

i dont understand this

OpenStudy (anonymous):

$$\Large{ a_n = a_1\cdot r^{n-1} \\~\\ a_1 = 4 ~, ~ a_8 =-8748 \\~\\\Rightarrow ~ a_n = 4\cdot r^{n-1} \\~\\\Rightarrow ~ a_8 = 4\cdot r^{8-1} \\~\\\Rightarrow ~ -8748 = 4\cdot r^{8-1} }$$

OpenStudy (anonymous):

what do I do after that? @jayzdd

OpenStudy (anonymous):

-2187 @jayzdd

OpenStudy (anonymous):

$$ \Large{ a_n = a_1\cdot r^{n-1} \\~\\ a_1 = 4 ~, ~ a_8 =-8748 \\~\\\Rightarrow ~ a_n = 4\cdot r^{n-1} \\~\\\Rightarrow ~ a_8 = 4\cdot r^{8-1} \\~\\\Rightarrow ~ -8748 = 4\cdot r^{8-1} \\~\\\Rightarrow ~ \frac{-8748}{4} = r^{8-1} \\~\\\Rightarrow ~ \frac{-8748}{4} = r^{7} \\~\\\Rightarrow ~ -2187 = r^{7} \\~\\\Rightarrow ~ (-2187)^{1/7} = r \\~\\\Rightarrow ~ -3= r } $$

OpenStudy (anonymous):

yeah then what do I do ?

OpenStudy (anonymous):

@jayzdd

OpenStudy (anonymous):

$$ \Large{ a_n = a_1\cdot r^{n-1} \\~\\ a_1 = 4 ~, ~ a_8 =-8748 \\~\\\Rightarrow ~ a_n = 4\cdot r^{n-1} \\~\\\Rightarrow ~ a_8 = 4\cdot r^{8-1} \\~\\\Rightarrow ~ -8748 = 4\cdot r^{8-1} \\~\\\Rightarrow ~ \frac{-8748}{4} = r^{8-1} \\~\\\Rightarrow ~ \frac{-8748}{4} = r^{7} \\~\\\Rightarrow ~ -2187 = r^{7} \\~\\\Rightarrow ~ (-2187)^{1/7} = r \\~\\\Rightarrow ~ -3= r \\~\\\Rightarrow ~a_n = 4\cdot (-3)^{n-1} \\~\\\Rightarrow ~a_{16} = 4\cdot (-3)^{16-1} } $$

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

@jayzdd

OpenStudy (anonymous):

yes thats correct

OpenStudy (anonymous):

Thank you so much @jayzdd

OpenStudy (anonymous):

Identify the 31st term of an arithmetic sequence where a1 = 26 and a22 = -226. (2 points) Question 2 options: 1) -334 2) -274 3) -284 4) -346

OpenStudy (anonymous):

can you help me with this one @jayzdd

OpenStudy (anonymous):

can you help me with this one @jayzdd